OpenStudy (anonymous):

-14*sqrt(x+5)*-14*sqrt(x+5) = 196*(x+5) right?

OpenStudy (btaylor):

yep

OpenStudy (anonymous):

So Distributive property applies to Radicals then right?

OpenStudy (btaylor):

yes. moreso, \[\sqrt a \times \sqrt b = \sqrt{ab}\]